The probability that Juan will win a game of chess whenever he plays is ¼. If he plays twice, what are the odds that he wins the 1st game and loses the 2nd game?
1
Expert's answer
2020-03-11T14:25:46-0400

Probability of winning the first game: 1/4;

Probability of losing the second game: 1-1/4=3/4

Probability of winning the first and losing the second game:

"\frac{1}{4}\cdot\frac{3}{4}=\frac{3}{16}"